OpenStudy (anonymous):

Create a quadratic polynomial function f(x) and a linear binomial in the form (x − a). Part 1. Show all work using long division to divide your polynomial by the binomial. Part 2. Show all work to evaluate f(a) using the function you created. Part 3. Use complete sentences to explain how the remainder theorem is used to determine whether your linear binomial is a factor of your polynomial function

OpenStudy (anonymous):

@ranga @annas @SolomonZelman

OpenStudy (solomonzelman):

@yeah_i_suck_at_math can you create a quadratic polynomial function and a linear polynomial as indicated in your question ?

OpenStudy (anonymous):

no

OpenStudy (solomonzelman):

Do you know what a "quadratic polynomial function" is? Do you know what a "linear binomial" is?

OpenStudy (anonymous):

no

OpenStudy (solomonzelman):

then I can't help you. Sorry -:(

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

@ranga

OpenStudy (anonymous):

Create one trinomial function: f(x) = x^3 - 3x^2 - x + 3., and a linear function g(x) = (x - 1). Part 1. Divide f(x) by g(x). You must do it yourself using the long division you have learned f(x)/g(x) = (x^3 - 3x^2 + x + 3)/(x - 1) = x^2 - 2x + 3 (no remainder) Part 2: a = 1 -> f(a) = f(1) = 0 Part 3. Use the remainder theorem?

OpenStudy (anonymous):

it would be nice if you expanded your answer but if you cannot , ill see what i can do